Computing EP: Stock Options Rand, Inc. had a net income of $80,000. During the year, 200,000 shares were outstanding on average and Rand’s common stock sold at an average market price of $50. In addition, Rand had 20,000 stock options outstanding to purchase a total of 20,000 common shares at $25 for each option exercised.

a. Compute basic earnings per share.

b. Compute diluted earnings per share.
Answer #1

a. Basic earnings per share

Since there is no mention of preference shares, it is safe to assume that there are no preference shares and therefore, no preference dividends.

Basic earnings per share can be found by dividing the net income available to equity shareholders(net income-preference dividend) by the weighted average number of equity shares outstanding during the year.

Basic earnings per share= 80000/200000= $0.4 per share

b. Diluted earnings per share

For computing the diluted eps we have to take into account the effect of potential equity shares also while computing the denominator ( outstanding shares). Besides this difference, the formula remains the same for diluted EPS also.

Amount paid if options are exercised= 20000*25= $500000

The above value in current shares= 500000/50= 10000 shares

Diluted shares= options available- value in current shares= 20000-10000= 10000 shares

Diluted EPS= (net income-preference dividend)/(outstanding shares+ diluted shares)= (80000-0)/(200000+10000)= $0.38 per share

Diluted EPS is generally lower than basic EPS.